Consider the following numerical example of the simple Keynesian model with no government spending, taxes or foreign sector(all figures in R million):
C=100+0,9Y
I=50
a, what is the value of the marginal prosperity to consume (MPC) in the model.

b, use a graph to illustrate the equilibrium level of output.
1
Expert's answer
2021-09-15T16:17:12-0400

Here , We consider a simple Keynesian Model with no government spending , taxes or foreign sector: 

The simple consumption function is of the form :

C = a + bY 

         where , C = Consumption spending 

                Y = Output / Income level 

                a = Autonomous consumption

                b = Marginal Propensity to Consume (MPC)

We are given   C = 100 + 0.9Y    

Therefore , the 

a.) MPC (Marginal Propensity to Consume): The MPC can be defined as the ratio of the additional increase/change in consumption due to the increase/ change in income. 

We are given   C = 100 + 0.9Y    

Therefore , the , MPC = 0.9 


b.) I = 50

We know at equilibrium, S= I 

                "S = 50\\ 

 Also , \space Y = C + S \\

 Y = 100 + 0.9Y + 50 \\

 Y - 0.9Y = 150\\

 0.1Y = 150 \\

 Y = 1500"

Therefore , the equilibrium level of output is 1500 R Millions